OpenStudy (anonymous):

how to solve this? find the cartesian equation of each of the curves with the following parametric equations: x=t^2 + t , y= t^2 - t

OpenStudy (anonymous):

you need to eliminate t from these two equations

OpenStudy (anonymous):

i`ve tried to eliminate t but stil cant get the answer..:(

OpenStudy (anonymous):

x+y=2 t^2 x-y= 2 t so, t= (x-y)/2 t^2= (x-y)^2/4 again, from equation 1, t^2= (x+y)/2 so, x+y=(x-y)^2/2

OpenStudy (anonymous):

good answer